# Real-time propagators resummed with nontrivial boundary wavefunctions in a constant electric field

### v1

@@ Abstract
We present the derivation of an alternative representation of the real-time in-in formalism under a spatially homogeneous and time independent electric field. Because the system exhibits instability associated with pair production of particles and antipartiles, the perturbation theory should be reorganized depending on the choice of the reference vacuum. We recast the boundary wavefunctions into the quadratic self-energy-like terms in the functional integration formalism. The resulting generating functional in the modified in-in formalism leads to the propagators that resum infinite diagrams necessary to capture the vacuum-instability effects. The proper-time representations of the propagators reproduce the known expressions from the canonical operator formalism, but our derivation based on the generating functional along the closed-time path clarifies the origin of the addtional proper-time contour and provides a better physical understanding. Finally, as a concrete example of the application, we compute the in-in expectation value of the vector current in a constant electric field, and find that the one-loop calculation results in the expression consistent with the physical intuition.
@@ 1 Introduction
The in-in (or the closed-time-path or the Schwinger-Keldysh) formalism is the standard real-time framework in quantum field theory Schwinger:1960qe ; Keldysh:1964ud ; Berges:2004yj . In practice, one often builds perturbation theory from the in-out propagators with the usual Feynman boundary conditions implemented by the $\mathrm{i}\epsilon$ prescription Weinberg:2005vy ; Calzetta:2008iqa . This choice is adequate for stable vacua, but in a background electric field it does not by itself capture the physics of vacuum decay. Observables directly tied to vacuum instability (e.g. pair production and the induced current) then require a consistent resummation of infinitely many diagrams Fedotov:2022ely . The issue is not only technical: when one tests operator relations such as Ward identities or energy-balance relations in real time, both sides must be resummed on the same footing Copinger:2018ftr , which makes a straightforward loop expansion cumbersome and can obscure the physical interpretation.
@@
In this work we provide such an embedding. The key observation is that, in a constant electric field, the definition of particles and vacuum is not unique: the in- and out-vacua are inequivalent and related by a Bogoliubov transformation. This fact implies that the “in-in boundary condition” admits distinct but equivalent representations. We formulate the in-in generating functional so as to keep the boundary wavefunctions explicitly, and then reorganize the perturbative expansion by choosing the out-state as the reference vacuum. In the functional integral, the boundary wavefunctions can be recast into quadratic, boundary-localized kernels that enter the action as self-energy–like terms. Importantly, this reorganization leaves the interaction vertices intact and modifies only the quadratic part, so that the resulting real-time propagators are obtained from the Dyson series dictated by the boundary condition.
This formulation has some advantages. First, it provides a practical framework in which vacuum-instability effects are incorporated at the propagator level, so in-in expectation values can be computed without tracking an explicit infinite resummation. Second, it clarifies the proper-time structure of the resummed propagators: while proper-time expressions in an electric field are known in canonical treatments, the appearance of additional proper-time contours may look mystical. By deriving the propagators directly from the closed-time-path generating functional, we identify the origin of these additional proper-time contours and give them a transparent interpretation in terms of the nontrivial boundary wavafunctions associated with the unstable vacuum.
@@
This paper is organized as follows. In Sec. 2 we review the standard real-time propagators in a constant electric field and pinpoint why the naïve in-in setup built from in-out propagators fails to capture vacuum-instability observables without resummation. In Sec. 3 we discuss the role of inequivalent vacua and reformulate the real-time expectation values with the effects of nontrivial boundary wavefunctions. In Sec. 4 we construct the corresponding generating functionals on the closed-time path extending from the out-state to the in-state. In Sec. 5 we derive the resulting resummed real-time propagators and their proper-time representations, elucidating the origin of the additional proper-time contours. In Sec. 6 we present the one-loop computation of the vector current as an application. Section 7 contains our conclusions and outlook.
For definiteness, we consider a spatially homogeneous and temporally constant electric field and ignore magnetic fields. We assume that interactions are adiabatically switched off in the asymptotic past and future. Throughout this paper we use natural units, $\hbar=1=c$, and adopt the metric convention $g_{\mu\nu}=\mathrm{diag}(+,-,-,-)$.
@@ 2 Real-time Propagators between Unstable Vacua
Here, the relative coordinate is z:=xyz:=x-y and a short-hand notation, ${\bm{z}}_{\perp}^{2}:=z_{1}^{2}+z_{2}^{2}$, is introduced. Throughout this work, the electric field is imposed along the $x^{3}$ axis with $E>0$. Thus, the vector potential is chosen to be
@@
where $E>0$. For the spatial components of the three vectors, we will use a sloppy notation of not distinguishing $z_{i}$ and $z^{i}$ as long as no confusion arises. We note that the first term in the exponential appears from the Schwinger phase, i.e.,
$\exp\biggl(\mathrm{i}e\int_{x}^{y}\mathrm{d}\ell\cdot A\biggr.)=\exp[\mathrm{i}\frac{eE}{2}(x^{0}+y^{0})z_{3}\biggr]$ (5)
@@
Here, it is important to make a remark that the proper-time contour, $\Gamma_{\mathrm{c}}$, extends from $s=0$ to $s=+\infty$ along the real axis as shown in Fig. 1. If this propagator in Eq. (1) is naïvely applied to perturbative calculations, the leading-order results may become different from what is expected from physical intuitions. For example, suppose that the current along the electric field, $\langle\hat{j}^{3}(x^{0})\rangle=\langle\bar{\psi}\gamma^{3}\psi\rangle$, is interested, the expected answer should be nonzero due to pair particle production and acceleration under the electric field. However, if one employs $S_{0}(x,y)$ in Eq. (1) to calculate $\langle\hat{j}^{3}(x^{0})\rangle$ at one-loop level, one would find $\langle\hat{j}^{3}(x^{0})\rangle=0$ Fradkin:1991zq ; Gavrilov:1998hw ; Copinger:2018ftr . Although it is not a preferable answer in physics, there is no apparent flaw in the mathematical formulation.
@@
In the in-in formalism Weinberg:2005vy ; Calzetta:2008iqa , one computes the expectation value of some time-dependent operator in the Heisenberg picture using the $\mathcal{C}_{+}$ path (upper path) and the $\mathcal{C}_{-}$ path (lower path) along the closed-time path in Fig. 2. Because the action takes different forms on the upper and the lower branches of the path, the in-in propagator, $S(x,y)$, have four independent components, namely, $S^{11}(x,y)$, $S^{22}(x,y)$, $S^{12}(x,y)$, and $S^{21}(x,y)$. To determine the propagators uniquely in general, one must specify the boundary conditions. In the path-integral language, the boundary condition and the temporal evolution correspond to the boundary wavefunction and the Feynman kernel, respectively. For the insertion of the in-vacuum wavefunction, as discussed in Refs. Weinberg:1995mt ; Weinberg:2005vy , one does not have to write down the wavefunction explicitly in the $\mathrm{i}\epsilon$ prescription. In the naïve implementation of the in-in formalism, the $4\times 4$ propagators are written as Berges:2004yj ; Weinberg:2005vy ; Calzetta:2008iqa :
@@
As already mentioned, the one-loop calculation in terms of the above propagators (6) leads to $\langle\hat{j}^{3}(x^{0})\rangle=0$ even in the in-in formalism. In other words, one must take some proper resummation of diagrams to reach a physically natural conclusion of $\langle\hat{j}^{3}(x^{0})\rangle\neq 0$. The resummation program is in principle feasible, but it would need delicate treatments. The following example could better explain this point. Suppose that the energy balance relation is interested, i.e.,
$\partial_{\mu}\hat{T}^{\mu 0}=F^{0\mu}\hat{j}_{\mu}\,$, (8)
@@
$\langle\hat{j}^{3}(x^{0})\rangle=\frac{(eE)^{2}x^{0}}{4\pi^{3}}\,\mathrm{e}^{-\pi m^{2}/(eE)}\,$, (10)
which finds the well-known formula under the name of the Schwinger mechanism. It would be useful to mention that Schwinger identified the imaginary part of the effective action involving the famous exponential factor, but did not directly compute time-growing quantities like Eq. (10). We will give a detailed derivation of this Schwinger formula later in Sec. 6.
@@ 4.1 Formal expressions of the generating functionals
In the standard formulation of the generating functional in functional integration, it is not necessary to convolute the wavefunctions under the assumption that only the ground state should be dominant in the infinite past and future. In the presence of time-dependent background fields, however, the in-state and the out-state are not equivalent, and the definition of the generating functional must retain the discrepancy between these two states.
@@
where $\phi_{+}(+\infty)$ is a short-hand notation for $\phi_{+}(x^{0}=+\infty,{\bm{x}})$ and we use similar notations in this paper. The boundary wavefunctions formally read:
@@ 4.2 Explicit forms of the boundary wavefunctions
up to an overall constant. In the absence of external fields, the mode function with a proper normalization is $f(t)=\mathrm{e}^{-\mathrm{i}\Omega t}/\sqrt{2\Omega}$ with some (dimensionless) frequency $\Omega$, from which we find the following relation: $2\mathrm{i}f\dot{f}=f/f^{*}$, from which we immediately derive $2\mathrm{i}(\dot{f}/{f})=1/|f|^{2}$. We demand that $f(t)$ under external fields should satisfy this relation (see the mode functions in Appendix A). Then, we can further simplify Eq. (25) as
@@ 4.3 Combined expressions of the generating functionals
We plug Eq. (30) into Eq. (16) for sQED and Eq. (34) into Eq. (18) for QED, respectively. Then, we obtain the full functional integration representations of the generating functionals. Here, we note that our expression in Eq. (30) is decomposed into two pieces with and without $\zeta_{\bm{p}}(\infty)$. The exponential factor without $\zeta_{\bm{p}}(\infty)$ is unrelated to the Bogoliubov transformation, simply leading to the free Fock vacuum in the $\mathrm{i}\epsilon$ prescription of the path integral Weinberg:1995mt ; Weinberg:2005vy . Keeping this in mind, we can make further simplification.
@@ 5.1 Formal relations of the real-time propagators
It should be noted that these are not the free propagators. We use the subscript “0” to indicate the propagators without the effects of the boundary wavefunctions, in the same sense as $S_{0}(x,y)$ in Eq. (1). Including the boundary wavefunctions in a form of the self energy, we can write down the $4\times 4$ real-time propagators, among which two are
@@
Similarly, for the Dirac field, we can write down the $4\times 4$ real-time propagators. The diagonal parts without the boundary wavefunctions are defined by
@@ 5.2 Outline of deriving the propagators
We solve Eqs. (41) and (43) using the mode functions summarized in Appendix A, and obtain the following $4\times 4$ propagators:
@@
In the same way, we can write down the $4\times 4$ propagators for the Dirac field as
@@ 5.3 Proper-time integrals for the real-time propagators in a constant electric field
For actual calculations, it is important to note that the integrands, $f(s;x,y)$ and $g(s;x,y)$, have singularities at $s=0$ and $s=-\mathrm{i}\pi/(eE)$. The former one at $s=0$ corresponds to the ultraviolet divergence; in fact, $f(s\ll 1/(eE);x,y)$ has no dependence on $eE$ except for the Schwinger phase and the singularity has nothing to do with the electric field. The latter one at $s=-\mathrm{i}\pi/(eE)$, on the other hand, appears from the real-time process associated with the electric field. One should treat these singularities in the principal-value prescription Schwinger:1960qe ; DiPiazza:2018ofz .
Although these forms of the real-time propagators along the modified proper-time paths are not well-known, they were first derived in canonical quantization several decades ago Nikishov:1969tt and used in some papers such as, e.g., Refs. Fradkin:1991zq ; Gavrilov:2007hq ; Gavrilov:2012jk . In this work, as we stated in the introduction, we aim to rederive these expressions within the framework of the functional integration formalism. In this way, it is now clearer which part takes care of resummation and how it is technically performed. For more technical details that are common to canonical quantization, see Appendix C.
@@ 6 Applications
In contrast, the second integral along $\Gamma^{<}-\Gamma^{>}$ contains the singularity in connection to the pair-production effect. Because the integration contour can be deformed in the complex plane as long as no singularity is crossed, we can make $\Gamma^{<}-\Gamma^{>}$ shrink around $s=-\mathrm{i}\pi/(eE)$. A straightforward calculation shows that the principal-value integral at $s=-\mathrm{i}\pi/(eE)$ along $\Gamma^{<}-\Gamma^{>}$ is unchanged by adding the contribution from the lower semicircular contour around it. This enables us to change $\Gamma^{<}-\Gamma^{>}$ into $\Gamma_{\mathrm{R}}^{\mathrm{a}}$ as illustrated in Fig. 7. If $z_{0}^{2}-z_{3}^{2}>0$, the integration contribution from the upper infinitesimal semicircle (and thus from $\Gamma_{\mathrm{R}}^{\mathrm{a}}$) vanishes, and this simplifies as
@@
where $\theta^{+}(z_{3}^{2}-z_{0}^{2}):=[\theta(z_{0}-0^{+})+\theta(-z_{0}-0^{+})]\theta(z_{3}^{2}-z_{0}^{2})$ to define its value at $z_{0}=0$. We see the necessity of some prescription from the derivative, e.g., $\partial^{3}\theta(z_{3}^{2}-z_{0}^{2})=2z_{3}\delta(z_{3}^{2}-z_{0}^{2})=2\mathrm{sgn}(z_{3})\bigl[\delta(z_{3}-z_{0})+\delta(z_{3}+z_{0})\bigr]$. This expression is subtle; if we perform the $z_{3}$ integration first and then take the $z_{0}\to 0$ limit (i.e., $x_{0}\to y_{0}$) next, the result should be zero. However, if we take the $z_{0}\to 0$ limit, we carefully need to define the sign function, and the result after the integration about $z_{3}$ is not necessarily zero. Thus, this subtlety about the definition of $\theta(z_{3}^{2}-z_{0}^{2})$ at $z_{0}=0$ should be specified, and this form of $\theta^{+}(z_{3}^{2}-z_{0}^{2})$ defines how to treat this subtlety at $z_{0}=0$, which is given in Ref. Fradkin:1991zq .

### v3

# arXiv:2512.19337v3[hep-ph] 26 Feb 2026

@@ Abstract
We present the derivation of an alternative representation of the real-time in-in formalism under a spatially homogeneous and time independent electric field. Because the system exhibits instability associated with pair production of particles and antiparticles, the perturbation theory should be reorganized depending on the choice of the reference vacuum. We recast the boundary wavefunctions into the quadratic self-energy-like terms in the functional integration formalism. The resulting generating functional in the modified in-in formalism leads to the propagators that resum infinite diagrams necessary to capture the vacuum-instability effects. The proper-time representations of the propagators reproduce the known expressions from the canonical operator formalism, but our derivation based on the generating functional along the closed-time path clarifies the origin of the additional proper-time contour and provides a better physical understanding. Finally, as a concrete example of the application, we compute the in-in expectation value of the vector current in a constant electric field, and find that the simple one-loop calculation captures the pair production effect.
@@ 1 Introduction
The in-in (or the closed-time-path or the Schwinger-Keldysh) formalism is the standard real-time framework in quantum field theory Schwinger:1960qe ; Keldysh:1964ud ; Berges:2004yj . In practice, one often builds perturbation theory from the in-out propagators with the usual Feynman boundary conditions implemented by the $\mathrm{i}\epsilon$ prescription Weinberg:2005vy ; Calzetta:2008iqa . This choice is adequate for stable vacua, but in a background electric field it does not by itself capture the physics of vacuum decay. Observables directly tied to vacuum instability (e.g. pair production and the induced current) then require a consistent resummation of infinitely many diagrams Fedotov:2022ely ; Copinger:2025ovz . The issue is not only technical: when one tests operator relations such as Ward identities or energy-balance relations in real time, both sides must be resummed on the same footing Copinger:2018ftr , which makes a straightforward loop expansion cumbersome and can obscure the physical interpretation.
@@
In this work we provide such an embedding. The key observation is that, in a constant electric field, the definition of particles and vacuum is not unique: the in- and out-vacua are inequivalent and related by a Bogoliubov transformation. This fact implies that the “in-in boundary condition” admits distinct but equally valid representations. We formulate the in-in generating functional so as to keep the boundary wavefunctions explicitly, and then reorganize the perturbative expansion by choosing the out-state as the reference vacuum. In the functional integral, the boundary wavefunctions can be recast into quadratic, boundary-localized kernels that enter the action as self-energy–like terms. Importantly, this reorganization leaves the interaction vertices intact and modifies only the quadratic part, so that the resulting real-time propagators are obtained from the Dyson series dictated by the boundary condition.
This formulation has some advantages. First, it provides a practical framework in which vacuum-instability effects are incorporated at the propagator level, so in-in expectation values can be computed without tracking an explicit infinite resummation. Second, it clarifies the proper-time structure of the resummed propagators: while proper-time expressions in an electric field are known in canonical treatments, the appearance of additional proper-time contours may look mystical. By deriving the propagators directly from the closed-time-path generating functional, we identify the origin of these additional proper-time contours and give them a transparent interpretation in terms of the nontrivial boundary wavefunctions associated with the unstable vacuum.
@@
This paper is organized as follows. In Sec. 2 we review the standard real-time propagators in a constant electric field and pinpoint why the naïve in-in setup built from in-out propagators fails to capture vacuum-instability observables without resummation. In Sec. 3 we discuss the role of inequivalent vacua and reformulate the real-time expectation values with the effects of nontrivial boundary wavefunctions. In Sec. 4 we construct the corresponding generating functionals on the closed-time path. In Sec. 5 we derive the resulting resummed real-time propagators and their proper-time representations, elucidating the origin of the additional proper-time contours. In Sec. 6 we present the one-loop computation of the vector current as an application. Section 7 contains our conclusions and outlook.
For definiteness, we consider a spatially homogeneous and temporally constant electric field and ignore magnetic fields. We assume that interactions are adiabatically switched off in the asymptotic past and future. Throughout this paper we use natural units, $\hbar=1=c$, and adopt the metric convention $g_{\mu\nu}=\mathrm{diag}(+,-,-,-)$. In addition, we use the following shorthand notations: $\int_{{\bm{p}}}:=\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}},\quad\int_{{\bm{x}}}:=\int\mathrm{d}^{3}x$.
@@ 2 Real-time propagators between unstable vacua
Here, the relative coordinate is z:=xyz:=x-y and a shorthand notation, ${\bm{z}}_{\perp}^{2}:=z_{1}^{2}+z_{2}^{2}$, is introduced. Throughout this work, the electric field is imposed along the $x^{3}$ axis with $E>0$. Thus, the vector potential is chosen to be
@@
where $E>0$. We note that the first term in the exponential appears from the Schwinger phase, i.e.,
$\exp\biggl(\mathrm{i}e\int_{x}^{y}\mathrm{d}\ell\cdot A\biggr.)=\exp[\mathrm{i}\frac{eE}{2}(x_{0}+y_{0})z_{3}\biggr]$ (5)
@@
Here, it is important to make a remark that the proper-time contour, $\Gamma_{\mathrm{c}}$, extends from $s=0$ to $s=+\infty$ along the real axis as shown in Fig. 1. If this propagator in Eq. (1) is naïvely applied to perturbative calculations, the leading-order results may become different from what is expected from physical intuitions. For example, consider the current along the electric field, $\langle\hat{j}^{3}(x^{0})\rangle=\langle\bar{\psi}\gamma^{3}\psi\rangle$. One expects it to be nonzero due to pair production and acceleration under the electric field. However, if one employs $S_{0}(x,y)$ in Eq. (1) to calculate $\langle\hat{j}^{3}(x^{0})\rangle$ at one-loop level, one would find $\langle\hat{j}^{3}(x^{0})\rangle=0$ Fradkin:1991zq ; Gavrilov:1998hw ; Copinger:2018ftr . Although it is not a preferable answer in physics, there is no apparent flaw in the mathematical formulation.
@@
In the in-in formalism Weinberg:2005vy ; Calzetta:2008iqa , one computes the expectation value of some time-dependent operator in the Heisenberg picture using the $\mathcal{C}_{+}$ path (upper path) and the $\mathcal{C}_{-}$ path (lower path) along the closed-time path in Fig. 2. Because the action takes different forms on the upper and the lower branches of the path, the in-in propagator, $S(x,y)$, has four independent components, namely, $S^{11}(x,y)$, $S^{22}(x,y)$, $S^{12}(x,y)$, and $S^{21}(x,y)$. To determine the propagators uniquely in general, one must specify the boundary conditions. In the path-integral language, the boundary condition and the temporal evolution correspond to the boundary wavefunction and the Feynman kernel, respectively. For the insertion of the in-vacuum wavefunction, as discussed in Refs. Weinberg:1995mt ; Weinberg:2005vy , one does not have to write down the wavefunction explicitly in the $\mathrm{i}\epsilon$ prescription. In the naïve implementation of the in-in formalism, the $2\times 2$ propagators are written as Berges:2004yj ; Weinberg:2005vy ; Calzetta:2008iqa :
@@
As already mentioned, the one-loop calculation in terms of the above propagators (6) leads to $\langle\hat{j}^{3}(x^{0})\rangle=0$ even in the in-in formalism. In other words, one must take some proper resummation of diagrams to reach a physically natural conclusion of $\langle\hat{j}^{3}(x^{0})\rangle\neq 0$. The resummation program is in principle feasible, but it would need delicate treatments. The following example could better explain this point. Suppose that one is interested in the energy balance relation, i.e.,
$\partial_{\mu}\hat{T}^{\mu 0}=F^{0\mu}\hat{j}_{\mu}\,.$ (8)
@@
$\langle\hat{j}^{3}(x^{0})\rangle=\frac{(eE)^{2}x^{0}}{2\pi^{3}}\,\mathrm{e}^{-\pi m^{2}/(eE)}\,$, (10)
which finds the well-known formula under the name of the Schwinger mechanism. We will give a detailed derivation of this formula later in Sec. 6.
@@ 4.1 Formal expressions of the generating functionals
In the standard functional-integration formulation of the in-in generating functional, the boundary wavefunction is a trivial Gaussian under the assumption that only the ground state should be dominant in the infinite past. By contrast, in the presence of time-dependent background fields, the in-state and the out-state are inequivalent, and the definition of the generating functional which we derive below retains the discrepancy between these two states through the insertion of operators $\hat{\mathscr{U}}_{\rm S/D}$ and $\hat{\mathscr{U}}^{\dagger}_{\rm S/D}$ at the boundary. Due to these operators, the boundary wavefunction becomes nontrivial (non-Gaussian), which induces the nonlocal self-energy-like terms in the generating functional and thereby modifies the propagators.
@@
where $\phi_{+}(+\infty)$ is a shorthand notation for $\phi_{+}(x^{0}=+\infty,{\bm{x}})$ and we use similar notations in this paper. The boundary wavefunctions formally read:
@@ 4.2 Explicit forms of the boundary wavefunctions
up to an overall constant. If we normalize the mode function as $f(t)=\mathrm{e}^{-\mathrm{i}\Omega t}/\sqrt{2\Omega}$ with a (dimensionless) frequency $\Omega$, we find the following relation: $2\mathrm{i}f\dot{f}=f/f^{*}$, from which we immediately derive $2\mathrm{i}(\dot{f}/{f})=1/|f|^{2}$. We demand that mode functions under external fields should satisfy this relation (see Appendix A). Then, we can further simplify Eq. (25) as
@@ 4.3 Combined expressions of the generating functionals
We plug Eq. (30) into Eq. (16) for sQED and Eq. (34) into Eq. (18) for QED, respectively. Then, we obtain the full functional integration representations of the generating functionals. Here, we note that our expression in Eq. (30) is decomposed into two pieces with and without $\zeta_{\bm{p}}(\infty)$. The exponential factor without $\zeta_{\bm{p}}(\infty)$ is unrelated to the Bogoliubov transformation, simply representing the free Fock vacuum. This factor ensures the convergence of the functional integration at the boundary and provides the $\mathrm{i}\epsilon$ term in the action Weinberg:1995mt ; Weinberg:2005vy . Similarly for the Dirac field, the terms independent of $\sigma_{\bm{p}}$ represent the free Fock vacuum, and this part serves only to provide the $\mathrm{i}\epsilon$ term. Keeping this in mind, we can make further simplification.
@@ 5.1 Formal relations of the real-time propagators
It should be noted that these are not the free propagators. We use the subscript “0” to indicate the propagators without the effects of the boundary wavefunctions, in the same sense as $S_{0}(x,y)$ in Eq. (1). Including the boundary wavefunctions in a form of the self energy, we can write down the $2\times 2$ real-time propagators, among which two are
@@
Similarly, for the Dirac field, we can write down the $2\times 2$ real-time propagators. The diagonal parts without the boundary wavefunctions are defined by
@@ 5.2 Outline of deriving the propagators
We solve Eqs. (41) and (42) using the mode functions summarized in Appendix A, and obtain the following $2\times 2$ propagators:
@@
In the same way, we can write down the $2\times 2$ propagators for the Dirac field as
@@ 5.3 Proper-time integrals for the real-time propagators in a constant electric field
For actual calculations, it is important to note that the integrands, $f(s;x,y)$ and $g(s;x,y)$, have singularities at $s=0$ and $s=-\mathrm{i}\pi/(eE)$. The former one at $s=0$ corresponds to the ultraviolet divergence; in fact, $f(s\ll 1/(eE);x,y)$ has no dependence on $eE$ except for the Schwinger phase and the singularity has nothing to do with the electric field. The latter one at $s=-\mathrm{i}\pi/(eE)$, on the other hand, appears from the real-time process associated with the electric field. One has to treat these singularities in the principal-value prescription Schwinger:1960qe ; DiPiazza:2018ofz .
Although these forms of the real-time propagators along the modified proper-time paths are not well-known, they were first derived in canonical quantization several decades ago Nikishov:1969tt and used in some papers such as, e.g., Refs. Fradkin:1991zq ; Gavrilov:2007hq ; Gavrilov:2012jk . In this work, as we stated in the introduction, our aim is to rederive these expressions within the framework of the functional integration formalism, and this aim has now been achieved. In this way, it is now clearer which part takes care of resummation and how it is technically performed. An interpretation of this resummation in terms of pair production is given in Refs. Copinger:2024pai ; Copinger:2025ovz . For further technical details that are common to canonical quantization, see Appendix C.
@@ 6 Applications
In contrast, the second integral along $\Gamma^{<}-\Gamma^{>}$ contains the singularity in connection to the pair-production effect. Because the integration contour can be deformed in the complex plane as long as no singularity is crossed, we can make $\Gamma^{<}-\Gamma^{>}$ shrink around $s=-\mathrm{i}\pi/(eE)$. A straightforward calculation shows that within $\Gamma^{<}-\Gamma^{>}$, the contribution from the contour segment parallel to the real axis is equivalent to that from a lower semicircular detour around the singularity. This enables us to change $\Gamma^{<}-\Gamma^{>}$ into $\Gamma_{\mathrm{R}}^{\mathrm{a}}$ as illustrated in Fig. 7. If $z_{0}^{2}-z_{3}^{2}>0$, the integration contribution from the upper infinitesimal semicircle vanishes, and the contribution from a part of the contour parallel to the real axis also disappears in the limit where the radius of the infinitesimal contour tends to zero. Thus, the remaining integral is simplified as
@@
where $\theta^{+}(z_{3}^{2}-z_{0}^{2}):=[\theta(z_{0}-0^{+})+\theta(-z_{0}-0^{+})]\theta(z_{3}^{2}-z_{0}^{2})$ to define its value at $z_{0}=0$. We see the necessity of some prescription from the derivative, e.g., $\partial^{3}\theta(z_{3}^{2}-z_{0}^{2})=2z_{3}\delta(z_{3}^{2}-z_{0}^{2})=2\mathrm{sgn}(z_{3})\bigl[\delta(z_{3}-z_{0})+\delta(z_{3}+z_{0})\bigr]$. This expression is subtle; if we perform the $z_{3}$ integration first and then take the $z_{0}\to 0$ limit (i.e., $x_{0}\to y_{0}$) next, the result should be zero. However, if we take the $z_{0}\to 0$ limit first, we carefully need to define the sign function, and the result after the integration about $z_{3}$ is not necessarily zero. Thus, this subtlety about the definition of $\theta(z_{3}^{2}-z_{0}^{2})$ at $z_{0}=0$ should be specified, and this form of $\theta^{+}(z_{3}^{2}-z_{0}^{2})$ defines how to treat this subtlety at $z_{0}=0$. This definition is given in Ref. Fradkin:1991zq .